The earth is 8 light minutes from the sun and the speed of light is 3x10 to the power of 8 m/s. Lets assume that the earths orbit is exactly circular. It's mass is 5.972x10 to the power of 24kg.
Using the values given above, determine the magnitude of the gravitational force that the sun exerts on the earth
Answer #1

distance between earth and sun

r = 8 * 60* 3* 10^8 = 1.44* 10^11 m

force

F = GMm / r^2 = 6.67* 10^-11* 5.98* 10^24* 5.972* 10^24 / (1.44* 10^11)^2

F = 1.154* 10^21 N

=====

Comment before rate in case any doubt, will reply for sure.. goodluck